OpenStudy (anonymous):

Solve the following system of equations. 3x-8y=7 -3x+13y=-17

OpenStudy (whpalmer4):

\[3x-8y=7\]\[-3x+13y=-17\] What would you get if you added both equations together? Do it like a big addition problem: 3x - 8y = 7 -3x + 13y = -17 --------------- ?

OpenStudy (anonymous):

@whpalmer4 y=-2?

OpenStudy (whpalmer4):

Yes! (Sorry, didn't notice you had responded) Now use that value of y in either of the original equations to find the value of x. What do you get?

OpenStudy (anonymous):

I lost my pencil, ive me a sec lol

OpenStudy (whpalmer4):

Good mental practice to try doing it in your head!

OpenStudy (anonymous):

then plug x and y into both equations to see if they work correct?

OpenStudy (anonymous):

I got it! :D x=-3 and y=-2

OpenStudy (whpalmer4):

Did you check them? The answers are correct. Good job!